OpenStudy (anonymous):

find indicated term 3rd, (a+b)^{6}

OpenStudy (anonymous):

so you are going to expand using the binomial theorem. The kth term of: \[(a+b)^n\] is : \[\left(\begin{matrix}n \\ k-1\end{matrix}\right)a^{n-(k-1)}b^{k-1}\] so in your problem, n = 6, and k = 3, so just plugging in numbers we get: \[\left(\begin{matrix}6 \\ 3-1\end{matrix}\right)a^{6-(3-1)}b^{3-1} = 15a^4b^2\]

OpenStudy (anonymous):

if you need anything explained in more detail just ask :)

OpenStudy (anonymous):

ok thank you:)